# Large Language Model Evaluation via Matrix Nuclear-Norm

### v1

@@ 1 Introduction
Large language models (LLMs), such as Gemini (Gemini et al., 2023), Llama (Touvron et al., 2023), and GPT-4 (GPT-4 Achiam et al., 2023), have demonstrated remarkable performance across a variety of natural language processing (NLP) tasks (Zhao et al., 2023). They are not only transforming the landscape of NLP (Saul et al., 2005; Liu et al., 2023; Sawada et al., 2023) but also bring beneficial impacts on computer vision (Lian et al., 2023a; Wang et al., 2024) and graph neural networks (Zhang et al., 2024; Chen et al., 2024b), achieving stellar results on various leaderboards. Despite these advancements, assessing a model’s ability to compress information remains a crucial research challenge (Delétang et al., 2023). Compression focuses on efficiently distilling essential information from vast training datasets while discarding redundant elements, showcasing a model’s ability to learn and recognize the underlying structure of the data (Wei et al., 2024). LLMs are expected to perform this form of compression during their training process (Zhao et al., 2023). Specifically, in the early stages of training, after random initialization, the representations produced from the data are often chaotic. However, as training progresses, these representations become more organized, allowing the model to filter out unnecessary information. Hence, assessing an LLM’s capacity for information compression is crucial for understanding its learning efficiency and representational power.
Current compression evaluation methods, such as Matrix Entropy introduced by Wei et al. (2024), measure information compression efficiency through processing models’ output representations on datasets. However, the reliance of Matrix Entropy on Singular Value Decomposition (SVD) (Kung et al., 1983; Zhang, 2015) leads to significant computational complexity, typically $O(n^{3})$, which limits its applicability in large-scale models.
@@
To validate the effectiveness of the Matrix Nuclear-Norm, we first conducted preliminary experiments on two language models of differing sizes. The results indicated a consistent decrease in Matrix Nuclear-Norm values as model size increased, signifying enhanced compression capabilities. Subsequently, we performed inference experiments on two widely used benchmark datasets, AlpacaEval (Dubois et al., 2024) and Chatbot Arena (Chiang et al., 2024), which cover a diverse range of language generation tasks. These benchmarks facilitate a comprehensive assessment of model inference performance. Our experimental findings confirm that the Matrix Nuclear-Norm accurately measures model compression capabilities and effectively ranks models based on performance, demonstrating its reliability and efficiency in practical applications. Our empirical investigations yield the following insights:
  1. 1.
    Proposal of the Matrix Nuclear-Norm: We present a new method that leverages the nuclear norm, successfully reducing the computational complexity associated with evaluating language models from $O(n^{3})$ to $O(n^{2})$. This reduction minimizes dependence on SVD, making the Matrix Nuclear-Norm a more efficient alternative to Matrix Entropy.
  2. 2.
    Extensive Experimental Validation: We validated the effectiveness of the Matrix Nuclear-Norm on language models of various sizes. Results indicate that this metric accurately assesses model compression capabilities, with values decreasing as model size increases, reflecting its robust evaluation capability.
  3. 3.
    Benchmark Testing and Ranking: We conducted inference tests on widely used benchmark datasets, AlpacaEval and Chatbot Arena, evaluating the inference performance of models across different sizes and ranking them based on the Matrix Nuclear-Norm. The results demonstrate that this metric can efficiently and accurately evaluate the inference performance of medium and small-scale models, highlighting its broad application potential in model performance assessment.
@@ 2 Related Work
Evaluating Large Language Models: Metrics and Innovations. The evaluation of large language models (LLMs) is rapidly evolving, driven by diverse tasks, datasets, and benchmarks (Wang et al., 2023; Celikyilmaz et al., 2020; Zheng et al., 2023; Chang et al., 2024). Effective evaluation is crucial for enhancing model performance and reliability. Traditional metrics, such as accuracy, F1 score (Sasaki, 2007), BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004), assess model outputs against annotated labels, while perplexity and cross-entropy loss focus on input data. Recent innovations, like Matrix Entropy (Wei et al., 2024), provide a means to evaluate information compression through activation matrix entropy, shedding light on internal representations. However, the computational intensity of Matrix Entropy, reliant on SVD, highlights the urgent requirement for developing more efficient evaluation methods in the context of LLMs.
Scaling Laws and Performance Metrics. Recent studies have consistently shown that as models and datasets expand, performance metrics such as loss improve predictably (Hestness et al., 2017; Kaplan et al., 2020; Clark et al., 2022). This trend is observed across various domains, including NLP and computer vision (Hoffmann et al., 2022; Bahri et al., 2024). Consequently, both industry and academia have prioritized the development of larger models (Wei et al., 2022; Touvron et al., 2023). Notably, scaling laws reveal a power-law relationship in performance enhancement. In addition to traditional metrics like loss and task accuracy (Zhang et al., 2022; Dey et al., 2023; Biderman et al., 2023), we introduce the Matrix Nuclear-Norm as a novel metric for assessing information compression and redundancy reduction in LLMs.
Insights from Information Theory for Large Language Models. Information theory has significantly enhanced our understanding of neural networks. The information bottleneck concept has been pivotal in elucidating supervised learning dynamics (Tishby et al., 2000). Recently, researchers have applied information-theoretic principles to self-supervised learning in vision tasks (Tan et al., 2023; Skean et al., 2023). Additionally, lossless compression techniques for LLMs, utilizing arithmetic coding, have been explored (Chen et al., 2024a; Valmeekam et al., 2023; Delétang et al., 2023). In this paper, we build on these foundations to introduce the Matrix Nuclear-Norm, an information-theoretic metric aimed at effectively evaluating the information compression and redundancy reduction capabilities of LLMs.
@@ 3 Preliminaries
This section presents the fundamental concepts and metrics used in our study to assess model performance, specifically focusing on discriminability, diversity, and the nuclear norm.
@@ 3.1 Discriminability Measurement: F-norm
Higher discriminability corresponds to lower prediction uncertainty in the response matrix $A$, quantified using Shannon Entropy (Shannon, 1948):
$H(A)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{C}A_{i,j}\log(A_{i,j}).$ (1)
In this equation, $B$ represents the number of samples, while $C$ denotes the number of distinct categories. Minimizing $H(A)$ results in maximal discriminability (Vu et al., 2019; Long et al., 2016), achieved when each row contains only one entry equal to 1 and all other entries equal to 0, indicating complete prediction certainty. Alternatively, we can focus on maximizing the Frobenius norm $\|A\|_{F}$:
@@
The Frobenius norm serves as a measure of the overall magnitude of the entries in $A$ and also reflects prediction certainty.
Theorem 1. Given a matrix $A$, $H(A)$ and $\|A\|_{F}$ are strictly inversely monotonic.
The proof is detailed in Sect A.4 of the Appendix. Thus, minimizing $H(A)$ is equivalent to maximizing $\|A\|_{F}$. The bounds for $\|A\|_{F}$ are:
$\sqrt{\frac{B}{C}}\leq\|A\|_{F}\leq\sqrt{B},$ (3)
where the upper and lower bounds correspond to maximum and minimum discriminability, respectively.
@@ 3.2 Diversity Measurement: Matrix Rank
Prediction diversity involves the number of unique categories in matrix $A$, denoted as $C_{p}(A)$. The expected value $E_{C}$ represents the average number of categories:
$E_{C}=E_{A\sim\mathcal{D}}\left(C_{p}(A)\right).$ (4)
In practice, $C_{p}(A)$ can be approximated by the rank of $A$ when $\|A\|_{F}$ is near its upper bound:
$C_{p}(A)\approx\operatorname{rank}(A).$ (5)
@@ 3.3 Nuclear Norm
Theorem 2. When $\|A\|_{F}\leq 1$, the convex envelope of $\operatorname{rank}(A)$ is the nuclear norm $\|A\|_{\star}$. The theorem is proved in Fazel (2002).
The nuclear norm $\|A\|_{\star}$, defined as the sum of singular values of $A$, has significant implications for assessing model performance. With $\|A\|_{F}\leq\sqrt{B}$, we have:
$\frac{1}{\sqrt{D}}\|A\|_{\star}\leq\|A\|_{F}\leq\|A\|_{\star}\leq\sqrt{D}\cdot\|A\|_{F},$ (6)
where $D=\min(B,C)$. Therefore, maximizing $\|A\|_{\star}$ ensures high diversity and discriminability.
The upper bound of $\|A\|_{\star}$ is given by:
@@ 4.2 Matrix Nuclear-Norm
Calculating the nuclear norm of matrix $A$ requires calculating its SVD, with a time complexity of $O\left(\min(B^{2}C,BC^{2})\right)$, simplifying to $O(n^{3})$, where $n$ represents the maximum of dimensions $B$ and $C$. While this computation is manageable for smaller dimensions, it becomes increasingly time-consuming as dimensions grow. Furthermore, SVD may not converge in some cases, it is thus important to develop more efficient approximations of singular values.
Since $A$ is typically sparse, with only a few non-zero responses in each category, its singular values can be approximated by aggregating these non-zero responses.
Theorem 3. When $\|A\|_{F}$ approaches the upper bound $\sqrt{B}$, the $j$-th largest singular value $\sigma_{j}$ can be approximated as:
$\sigma_{j}\approx\text{top}\left(\sum_{i=1}^{B}A_{i,j}^{2},j\right),\quad\forall j\in\{1,\ldots,D\}.$ (8)
The proof is provided in Sect. A.5 of the Supplementary Materials. Using this approximation, the batch nuclear norm can be efficiently calculated as:
$\|\hat{A}\|_{*}=\sum_{j=1}^{D}\operatorname{top}\left(\sqrt{\sum_{i=1}^{B}A_{i,j}^{2}},j\right).$ (9)
This approach indicates that the primary components of the $L_{1,2}$-norm can effectively approximate the nuclear norm when $\|A\|_{F}$ is close to $\sqrt{B}$, while other components can be considered noise. Compared to traditional SVD-based methods, such as Guo et al. (2015), our Matrix Nuclear-Norm calculation offers two main advantages: (1) it reduces computational complexity from $O(n^{3})$ to $O(n^{2})$, and (2) it relies solely on ordinary floating-point calculations, avoiding risks of non-convergence. The complete algorithm for Matrix Nuclear-Norm is detailed in Algorithm 1.
@@
0:  Sentence representations (hidden states from LLM) $\mathcal{S}=\left\{X_{i}\right\}_{i=1}^{m}$, where $x_{i}\in\mathbb{R}^{d\times 1}$, $d$ is the hidden dimension of representation, and $L_{\text{input}}$ is the length of the sentence.
1:  $\mu=\frac{1}{m}\sum_{i=1}^{m}x_{i}$ // Calculate the mean embedding
2:  $\mathbf{X}_{\text{norm}}=\frac{\mathbf{X_{i}}-\mu}{\|\mathbf{X_{i}}-\mu\|_{2,\text{row}}}$ // Normalize the activation matrix
@@
[…]
@@ 5 Experiments of Large Language Models
5.1 Implementation Details
5.1.1 Baselines
Cross-Entropy Loss. For a text sequence $U=\{u_{1},\ldots,u_{T}\}$, the cross-entropy loss is defined as follows (Yang, 2019):
$L(U)=-\frac{1}{T}\sum_{i=1}^{T}\log P(u_{i}|u_{<i};\Theta)$ (11)
This loss quantifies the model’s ability to predict $u_{i}$ based on preceding words $u_{<i}$. We evaluate the performance of the Matrix Entropy metric in comparison to this baseline, utilizing the same datasets and models as referenced in (Kaplan et al., 2020).
Perplexity. The perplexity of a language model for a text sequence $U=\{u_{1},\ldots,u_{T}\}$ is defined as (Neubig, 2017):
$PPL(U)=\exp\left(-\frac{1}{T}\sum_{i=1}^{T}\log P(u_{i}|u_{<i};\Theta)\right)$ (12)
Perplexity measures the model’s capacity to predict the next word, reflecting the average number of attempts required for correct predictions.
Matrix Entropy of a Dataset. Given a dataset $\mathcal{D}=\{\mathcal{S}_{i}\}_{i=1}^{n}$, where $\mathcal{S}_{i}$ represents the samples (sentence embeddings) within the dataset, the (normalized) matrix entropy of $\mathcal{D}$ is defined as (Wei et al., 2024):
$H(\mathcal{D})=\frac{\sum_{i=1}^{n}H(\Sigma_{\mathcal{S}_{i}})}{n\log d}.$ (13)

### v3

@@ Abstract
Large Language Model Evaluation via Matrix Nuclear-Norm
Yahan Li1, Tingyu Xia1, Yi Chang1,2,3, Yuan Wu1,2thanks: Corresponding author 1School of Artificial Intelligence, Jilin University 2Key Laboratory of Symbolic Computation and Knowledge Engineering, Jilin University 3International Center of Future Science, Jilin University yahan23@mails.jlu.edu.cn, xiaty21@mails.jlu.edu.cn, yichang@jlu.edu.cn yuanwu@jlu.edu.cn
@@ 1 Introduction
Large language models (LLMs), such as Gemini (Gemini et al., 2023), Deepseek (Guo et al., 2025), and GPT-4 (GPT-4 Achiam et al., 2023), have shown exceptional performance in numerous natural language processing (NLP) tasks (Zhao et al., 2023). These models are transforming the way we approach NLP tasks, providing unprecedented capabilities and solutions to complex problems. They are revolutionizing NLP (Saul et al., 2005; Liu et al., 2023; Sawada et al., 2023) and positively impacting computer vision (Lian et al., 2023a; Wang et al., 2024) and graph neural networks (Zhang et al., 2024; Chen et al., 2024), achieving top results on leaderboards. Despite these advancements, evaluating a model’s ability to compress information remains a critical research challenge (Delétang et al., 2023). This challenge is essential for improving the overall efficiency of these models.
Compression involves efficiently extracting essential information from large datasets while removing redundant data, highlighting a model’s ability to understand the data’s underlying structure (Wei et al., 2024). LLMs are expected to perform this compression during training (Zhao et al., 2023). Initially, after random initialization, the data representations are chaotic, but as training progresses, they become organized, allowing the model to filter out unnecessary information. Thus, assessing an LLM’s compression capacity is vital for understanding its learning efficiency and representational power, which are crucial for practical applications and real-world deployment.
@@
To validate the Matrix Nuclear-Norm, we conducted preliminary experiments on two language models of different sizes. Results showed a consistent decrease in Matrix Nuclear-Norm values as model size increased, indicating enhanced compression capabilities. We also performed inference experiments on benchmark datasets, AlpacaEval (Dubois et al., 2024) and Chatbot Arena (Chiang et al., 2024), covering diverse language generation tasks. These benchmarks provide a comprehensive assessment of model inference performance. Our findings confirm that the Matrix Nuclear-Norm accurately measures model compression capabilities and ranks models based on performance, demonstrating its reliability and efficiency. Our empirical investigations yield the following insights:
  • Proposal of the Matrix Nuclear-Norm: We introduce a method leveraging the nuclear norm, reducing computational complexity from $O(n^{3})$ to $O(n^{2})$. This reduction minimizes SVD dependence, making Matrix Nuclear-Norm a more efficient alternative to Matrix Entropy.
  • Extensive Experimental Validation: We validated the Matrix Nuclear-Norm on language models of various sizes. Results show this metric accurately assesses model compression capabilities, with values decreasing as model size increases, reflecting its robust evaluation capability.
  • Benchmark Testing and Ranking: We conducted inference tests on benchmark datasets, AlpacaEval and Chatbot Arena, evaluating inference performance across different model sizes and ranking them based on the Matrix Nuclear-Norm. Results demonstrate this metric efficiently and accurately evaluates medium and small-scale models, highlighting its broad application potential in model performance assessment.
@@ 2 Related Work
LLM Evaluation and Scaling Laws. Evaluating large language models (LLMs) is a multifaceted challenge, as it requires capturing both task-specific performance and internal representational efficiency. Scaling laws have become a foundational framework for studying how LLM performance evolves with model size and data volume (Kaplan et al., 2020; Ruan et al., 2024). These studies demonstrate that model performance on tasks like language modeling and fine-tuning often follows predictable power-law relationships with respect to model parameters and dataset size, emphasizing the importance of scaling for achieving state-of-the-art results.However, scaling laws typically focus on external metrics such as cross-entropy loss, offering limited insight into how LLMs manage internal knowledge representation. For instance, the ability of LLMs to compress knowledge, eliminate redundancy, and retain structured information remains poorly understood with traditional methods. Addressing these gaps requires structural metrics that go beyond task outcomes to directly evaluate the internal embeddings and activation patterns of LLMs.
LLM Evaluation Metrics. Traditional evaluation metrics such as perplexity, BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004) primarily measure task-specific outcomes, assessing how well model outputs align with ground truth data. While these metrics are effective for evaluating surface-level outputs, they do not capture the underlying mechanisms of LLMs, such as the diversity or compression of embeddings. Similarly, accuracy and F1 score (Sasaki, 2007) focus on classification performance, making them less applicable to the generative tasks typical of LLMs.To bridge this gap, structural metrics such as Matrix Entropy have been introduced. Matrix Entropy (Wei et al., 2024) employs information theory to assess the entropy of covariance matrices derived from LLM embeddings. This metric evaluates how effectively a model removes redundancy and encodes structured information, offering a measure of its compression capabilities. For instance, Matrix Entropy can reveal differences in embedding distributions across models of varying sizes, reflecting their capacity to extract meaningful patterns from large datasets. However, its reliance on Singular Value Decomposition (SVD) results in a computational complexity of $O(n^{3})$, limiting its applicability to modern large-scale models. To overcome these limitations, we propose the Matrix Nuclear-Norm as a scalable alternative. By leveraging the $L_{1,2}$ norm as a convex approximation of matrix rank, the Matrix Nuclear-Norm reduces computational complexity to $O(n^{2})$. This makes it feasible for evaluating embeddings from large-scale LLMs while preserving the insights provided by Matrix Entropy, such as compression efficiency.
@@ 3 Preliminaries
This section presents the fundamental concepts for model performance evaluation: discriminability, diversity, and nuclear norm.
@@ 3.1 Discriminability Measurement: F-NORM
Higher discriminability corresponds to lower prediction uncertainty in the response matrix $A$. When $A$ is normalized as a probability matrix (i.e., $\sum_{j=1}^{C}A_{i,j}=1,\ \forall i\in[B]$), this uncertainty can be quantified using Shannon Entropy (Shannon, 1948):
$H(A)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{C}A_{i,j}\log\left(A_{i,j}\right)$ (1)
where $B$ is the number of samples, $C$ the feature dimension, and $A_{i,j}$ the normalized activation value. Lower entropy indicates higher discriminability.
An alternative measurement is the Frobenius norm:
@@
This norm reflects activation intensity, with higher values indicating more concentrated distributions.
Theorem 1. For a row-normalized matrix $A\in\mathbb{R}_{+}^{B\times C}$ (i.e., $\sum_{j=1}^{C}A_{i,j}=1,\ \forall i$), $H(A)$ and $\|A\|_{F}$ are strictly inversely monotonic.
The norm satisfies dimensional bounds:
$\sqrt{\frac{B}{C}}\leq\|A\|_{F}\leq\sqrt{B}$ (3)
where the lower bound achieves when $A$ has uniform distributions (maximal uncertainty), and the upper bound when $A$ contains one-hot vectors (minimal uncertainty). The proof is given in Appendix A.5.
@@ 3.2 Diversity Measurement: Matrix Rank
In LLMs, diversity reflects the model’s ability to utilize its latent representation space effectively. For a given dataset $\mathcal{D}$, the expected diversity of outputs is defined as:
$E_{C}=\mathbb{E}_{A\sim\mathcal{D}}\big{[}C_{p}(A)\big{]}$ (4)
To approximate $C_{p}(A)$, we construct a sparse matrix $M\in\{0,1\}^{B\times C}$ where each row contains a one-hot vector indicating the argmax position:
$M_{i,j}=\begin{cases}1,&j=\arg\max_{k}A_{i,k}\\ 0,&\text{otherwise}\end{cases}$ (5)
The capacity measure then becomes:
$C_{p}(A)=\mathrm{rank}\big{(}M\odot A\big{)}\approx\mathrm{rank}(A)$ (6)
where $\odot$ denotes element-wise product.
The maximum value of $C_{p}(A)$ is $\min(B,C)$, where $C$ is the output representation dimension. Maximizing $C_{p}(A)$ ensures effective utilization of the representation space, promoting robustness through reduced redundancy.
@@ 3.3 Nuclear Norm
Theorem 2. When $\|A\|\leq 1$ (where $\|A\|$ is the spectral norm), the convex envelope of $\operatorname{rank}(A)$ is the nuclear norm $\|A\|_{\star}$. The theorem is proved in Fazel (2002).
For a matrix $A\in\mathbb{R}^{B\times C}$ with $\|A\|_{F}\leq\sqrt{B}$, let $D=\min(B,C)$. The relationships between the nuclear norm and Frobenius norm are:
$\|A\|_{F}\leq\|A\|_{\star}\leq\sqrt{D}\cdot\|A\|_{F}.$ (7)
Therefore, maximizing $\|A\|_{\star}$ encourages higher rank, which implies high diversity and discriminability. The upper bound of $\|A\|_{\star}$ is further bounded by:
@@ 4.2 Matrix Nuclear-Norm
For a matrix $A\in\mathbb{R}^{B\times C}$, computing its exact nuclear norm via Singular Value Decomposition (SVD) requires $O(\min(B^{2}C,BC^{2}))$ time, which is equivalent to $O(n^{3})$ with $n=\max(B,C)$. While feasible for small matrices, this becomes computationally prohibitive for large-scale models. Additionally, numerical instability may arise in SVD computations for ill-conditioned matrices.
Sparsity Prior: When $A$ exhibits column-wise sparsity (i.e., non-zero activations concentrate in a subset of columns), we can approximate its singular values by leveraging column norms. Let $\|A\|_{F}$ denote the Frobenius norm, bounded by $\|A\|_{F}\leq\sqrt{\min(B,C)}\cdot\sigma_{\max}(A)$, where $\sigma_{\max}(A)$ is the largest singular value.
Theorem 3. (Column-Norm Approximation)
If $A$ has rapidly decaying column norms $\{\|A_{:,j}\|_{2}\}_{j=1}^{C}$, the $j$-th largest singular value $\sigma_{j}(A)$ can be approximated by the $j$-th largest column norm:
$\sigma_{j}(A)\approx\mathrm{Sort}\left(\left\{\|A_{:,j}\|_{2}\right\}_{j=1}^{C}\right)_{[j]},\quad j\in\{1,\dots,r\},$ (9)
where $r=\mathrm{rank}(A)$. The proof is given in Sect. A.6 (Supplementary Materials). The nuclear norm is then approximated as:
$\|\hat{A}\|_{\star}\approx\sum_{j=1}^{D}\mathrm{Sort}\left(\left\{\|A_{:,j}\|_{2}\right\}_{j=1}^{C}\right)_{[j]},$ (10)
where $D\leq r$ is a hyperparameter controlling approximation precision, and $\widetilde{A}$ denotes the column-sparse approximation of $A$.
Remark: This approximation holds under the assumption that off-diagonal correlations between columns are negligible (i.e., $A^{\top}A\approx\mathrm{diag}(\|A_{:,1}\|_{2}^{2},\dots,\|A_{:,C}\|_{2}^{2})$). For correlated columns, a diagonal correction term may be required.
This approach indicates that the primary components of the $L_{1,2}$-norm can effectively approximate the nuclear norm when $\|A\|_{F}$ is close to $\sqrt{B}$, while other components can be considered noise. Compared to traditional SVD-based methods (e.g., Guo et al. (2015)), this approach reduces computational complexity from $O(n^{3})$ to $O(n^{2})$ and avoids convergence issues by using only standard floating-point operations. The complete algorithm is detailed in Algorithm 1.
@@
0:  Sentence representations $\mathcal{S}=\left\{X_{i}\right\}_{i=1}^{m}$, where $X_{i}\in\mathbb{R}^{d\times 1}$, $d$ is the hidden dimension, and $L_{\text{input}}$ is the sentence length.
1:  $\mu=\frac{1}{m}\sum_{i=1}^{m}X_{i}$ // Mean embedding
2:  $\mathbf{X}_{\text{norm}}=\frac{\mathbf{X}-\mu}{\|\mathbf{X}-\mu\|_{2,\text{row}}}$ // Normalize matrix
@@
4:  $\Sigma_{D}=\{\sigma_{1},\sigma_{2},\dots,\sigma_{D}\}$ // Top $D$ norms
@@ 5 Experiments of Large Language Models
The models and datasets used in this paper are thoroughly introduced in  A.2.
5.1 Baselines
Cross-Entropy Loss. Cross-entropy is a key metric for evaluating LLMs by measuring the divergence between predicted and true probability distributions. The formula is given as (Wei et al., 2024):
$\mathcal{L}_{\text{CE}}=-\frac{1}{T}\sum_{i=1}^{T}\log P(u_{i}\mid u_{<i};\Theta)$ (12)
where $u_{i}$ is the target token at position $i$, $P(u_{i}\mid u_{<i};\Theta)$ is the conditional probability predicted by the model, and $T$ is the sequence length. Lower values indicate better prediction accuracy. We compare this baseline with the Matrix Nuclear Norm metric, using the same datasets and models from (Kaplan et al., 2020).
Perplexity. Perplexity measures how well a language model predicts a sequence of words. For a text sequence $\mathbf{U}=\{u_{1},\ldots,u_{T}\}$, it is defined as (Neubig, 2017; Wei et al., 2024):
$\text{PPL}(\mathbf{U})=\exp\left(-\frac{1}{T}\sum_{i=1}^{T}\log P(u_{i}\mid u_{<i};\Theta)\right)$ (13)
Lower perplexity indicates better performance, showing that fewer attempts are needed to predict the next token.
Matrix Entropy of a Dataset. For a dataset $\mathcal{D}=\{\mathbf{S}_{i}\}_{i=1}^{n}$, where $\mathbf{S}_{i}\in\mathbb{R}^{d\times d}$ represents sentence embedding covariance matrices, the normalized matrix entropy is defined as (Wei et al., 2024):
$H(\mathcal{D})=\frac{1}{n\log d}\sum_{i=1}^{n}H\left(\frac{\sigma(\mathbf{S}_{i})}{\|\sigma(\mathbf{S}_{i})\|_{1}}\right)$ (14)
where $\sigma(\mathbf{S}_{i})$ denotes the singular values of matrix $\mathbf{S}_{i}$, and $H(\cdot)$ is the Shannon entropy computed over the normalized singular value distribution.